OpenStudy (anonymous):

Metal will be giving ;) Which is the equation of the line? a.4x-7y=28 b.4x+7y=-28 c.-4x+7y=28 d.-4x-7y=-28

undeadknight26 (undeadknight26):

What is the given line?

OpenStudy (anonymous):

I know the answer but can you atleast try to plot it.

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

To find x intercept : To find y: Ax + By = C Ax+By=C 4x-7y=28 4x-7y=28 2(0) -7y=28 4x-7(0)=28 2*0=0 7*0=0 -7y=28 4x=28 28÷-7=-4 28/4 =7 x = -4 y=7

OpenStudy (anonymous):

Does that answer your question or do you need further explanation.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so the answer would be A.